OpenStudy (anonymous):

x^4-9x^2+c=0 a) Find the value of c such that the equation has four distinct real roots b) Find the value of c such that the equation has two distinct roots c) Find the value of c such that the equation has no roots

OpenStudy (anonymous):

how do I begin?

OpenStudy (asnaseer):

HINT: replace \(x^2\) with \(y\)

OpenStudy (anonymous):

ok.then y^2-9y+c=0. then..

OpenStudy (asnaseer):

do you know about the "discriminant" of a quadratic equation?

OpenStudy (anonymous):

No.

OpenStudy (raden):

let x^2 = y the equation will be : y^2 - 9y + c = 0 then take D > 0

OpenStudy (asnaseer):

this might help you learn about this property: http://www.mathwarehouse.com/quadratic/discriminant-in-quadratic-equation.php

OpenStudy (anonymous):

is it use b^2-4ac.

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

the value of the discriminant determines the number of real roots that a quadratic equation has

OpenStudy (asnaseer):

if we call the discriminant D, then: D = 0 => two equal real roots D > 0 => two distinct real roots D < 0 => no real roots

OpenStudy (anonymous):

C > 81/4.but...how to get 4 roots

OpenStudy (asnaseer):

remember you substituted \(y=x^2\) to get:\[y^2-9y+c=0\]so if you get two real roots for y, then that means 4 real roots for x since \(x=\pm\sqrt{y}\)

OpenStudy (anonymous):

I still didn't understand. is my ans correct?

OpenStudy (asnaseer):

do you agree that C > 81/4 will give you two rel roots for y?

OpenStudy (anonymous):

No. how to solve....pls step by step

OpenStudy (asnaseer):

you need to understand the significance of discriminant fully in order to understand how to solve this. I suggest you first spend some time understanding that and then come back to this problem. Use the web link I gave above or just search for "quadratic discriminant" in google.

OpenStudy (asnaseer):

if you are still stuck afterwards then give me a ping and I will try to explain further.

OpenStudy (asnaseer):

ok - so do you understand the discriminant fully now?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

good, so what would be the discriminant for:\[y^2-9y+c=0\]

OpenStudy (anonymous):

D>0

OpenStudy (asnaseer):

no - what is the expression for the discriminant for that equation?

OpenStudy (asnaseer):

e.g. if we have: \(ax^2+bx+c=0\) then the discriminant would be equal to \(b^2-4ac\)

OpenStudy (anonymous):

81-4c>0

OpenStudy (asnaseer):

there is no ">0" here. the discriminant is equal to 81-4c agreed?

OpenStudy (anonymous):

agree

OpenStudy (asnaseer):

good - so if now call the discriminant D, then we can write:\[D=81-4c\]now what are the 3 rules for the value of the discriminant?

OpenStudy (anonymous):

as u ve mentioned bf, D = 0 => two equal real roots D > 0 => two distinct real roots D < 0 => no real roots

OpenStudy (asnaseer):

perfect - so now find what value(s) of 'c' will lead to each of the 3 above cases. what do you get?

OpenStudy (anonymous):

81-4c>0 81-4c<0 81-4c=0

OpenStudy (anonymous):

is that so?

OpenStudy (asnaseer):

yes, and now solve each of these to find 'c' in each case

OpenStudy (anonymous):

c>81/4 c<81/4 c=81/4

OpenStudy (asnaseer):

1st two are incorrect

OpenStudy (anonymous):

how do I solve then?

OpenStudy (asnaseer):

please list the steps you took to solve the 1st one and I will try to spot where you may have made a mistake

OpenStudy (anonymous):

81-4c>0 81-4c<0 is it c < 81/4 c> 81/4......bit confused

OpenStudy (anonymous):

>=?

OpenStudy (anonymous):

where did I go?

OpenStudy (asnaseer):

take a simpler case: 81-c>0 what does this mean for 'c'?

OpenStudy (anonymous):

c<81

OpenStudy (asnaseer):

exactly - no retry the case we had above please

OpenStudy (asnaseer):

*now

OpenStudy (anonymous):

c<81/4 c>81/4 c=81/4

OpenStudy (asnaseer):

perfect, so now you know that: c = 81/4 => two equal real roots for y c < 81/4 => two distinct real roots for y c > 81/4 => no real roots for y agreed?

OpenStudy (anonymous):

after solving for c, I ve to subst back to the eqn. as \[y^2-9y+c=0\]

OpenStudy (asnaseer):

you don't need to substitute back into the equation

OpenStudy (anonymous):

\[y^2-9y+81/4=0\]

OpenStudy (asnaseer):

you are not asked to find the roots - you are asked to find the value for 'c' to give you a certain number of real roots

OpenStudy (anonymous):

Sry, hw to determine the roots?

OpenStudy (asnaseer):

do you agree that you have been asked to find the values for 'c' that give you: a) four distinct real roots b) two distinct real roots c) no real roots for the equation: \(x^4-9x^2+c=0\)

OpenStudy (anonymous):

Yes. surely

OpenStudy (asnaseer):

so, we first substituted \(y=x^2\) to get: \(y^2-9y+c=0\) and found that: a) c < 81/4 => two distinct real roots for y b) c = 81/4 => two equal real roots for y c) c > 81/4 => no real roots for y agreed?

OpenStudy (anonymous):

agree.

OpenStudy (asnaseer):

so, finally, we know that \(x=\pm\sqrt{y}\) which means: a) two distinct real roots for y => four distinct real roots for x b) two equal real roots for y => two distinct real roots for x c) no real roots for y => no real roots for x

OpenStudy (asnaseer):

remember, for each root of y we get two roots for x since \(x=\pm\sqrt{y}\)

OpenStudy (asnaseer):

we can therefore say that: a) c < 81/4 => four distinct real roots for x b) c = 81/4 => two distinct real roots for x c) c > 81/4 => no real roots for x

OpenStudy (anonymous):

OK. thanks a lot!!!!! Clearly explained. Cn I ask a qn on inequality?

OpenStudy (asnaseer):

<<<---- sure - just post it as a new question on the left